OpenStudy (anonymous):

The function f: R--->R defined as \[f(x) =\frac{ 3x^2 + 3x - 4 }{ -4x^2 + 3x + 3 }\] check wheather it is one one or on to funtions

OpenStudy (zarkon):

have you graphed it?

OpenStudy (anonymous):

Nope..)

OpenStudy (anonymous):

(4y+3)x2+(3-3y)x-(4+3y)=0 check delta if delta>0 it is not one to one

OpenStudy (anonymous):

if for all y delta<=0 so it is one to one

OpenStudy (anonymous):

delta=b2-4ac

OpenStudy (anonymous):

because for a y it gives you two x for some y.

OpenStudy (anonymous):

delta=(3-y)2+(4+3y)2 >0 so it is not one to one.

OpenStudy (anonymous):

yup....@mahmit2012 that make lot of sense....then hw to check for onto condition..?

OpenStudy (anonymous):

for all y#-3/4 you have some x for y=-3/4 check .you have also x=0

OpenStudy (anonymous):

so it is onto also

OpenStudy (anonymous):

not zero but there is a x

OpenStudy (sirm3d):

BUT the codomain is R, not R\{-3/4} so it is NOT onto since y=-3/4 has no pre-image x.

OpenStudy (anonymous):

why I checked y=-3/4 because it is x2 coefficent or horizental asymtotic line

OpenStudy (anonymous):

@Zarkon right, but for your stage and second order equation that's better to figure out with delta.

OpenStudy (sirm3d):

sorry, my mistake.

OpenStudy (sirm3d):

but strictly speaking, the relation is not a function since the relation is not well defined, i.e., there are two values of x that makes the denominator zero.

OpenStudy (anonymous):

y=-3/4-> 21x/9=-7/4 x=-3/4 check it out!

OpenStudy (anonymous):

if both coefficient of x and x2 were the same so you right, but it is not @sirm3d

OpenStudy (zarkon):

the question states that "f: R--->R" which is not true

OpenStudy (anonymous):

@mahmit2012 i did nt get u ( onto) ?

OpenStudy (anonymous):

onto means for all y you have at least one x. so the second equation tells you it is.

OpenStudy (anonymous):

wat is second eq here

OpenStudy (anonymous):

@Zarkon That is related to domain not for function to be onto.

OpenStudy (anonymous):

f:A->B for checking onto you should pay attention to B and cover B in Domain belonging to A.

OpenStudy (zarkon):

the function is ill defined if they write "f:R->R"

OpenStudy (zarkon):

they should write f:A->R where A is the maximal set in which f can be defined